# Sustainable Volunteer Engagement: Ensuring Potential Retention and Skill Diversity for Balanced Workforce Composition in Crowdsourcing Paradigm

### v1

@@ Abstract.
ccs: Human-centered computing Collaborative and social computing theory, concepts and paradigmsccs: Information systems Crowdsourcing
1. Introduction
Crowdsourcing (CS) has become integral to gathering and analyzing diverse data for various social applications like volunteering, delivery logistics, and participatory sensing (Bedogni et al., 2023). These activities involve diverse crowd workers who contribute actively or passively, dealing with spatial or non-spatial data in various formats and qualities.
@@ 1. Introduction
Research shows that most of the crowd workers on online CS platforms are primarily short-term participants. This deficiency in sustained crowd engagement has led to an imbalance in workforce composition, hindering large market enterprises from achieving their full utility (Shi et al., 2022). Crowd workers often disengage due to waning motivation or switch platforms for better perks and visibility, highlighting the need for strategies to select and retain competent workers (Nawaz et al., 2019). While initial participation is vital in CS platforms, sustained crowd engagement is crucial for their long-term success and the continuous delivery of high-quality services for social good (Leung et al., 2021; Wu et al., 2022).
@@
Volunteer Crowdsourcing Services (VCS), such as VolunteerMatch (VolunteerMatch, 2021) and SkillImpact (Skilled Impact, 2021), are platforms within the growing community of citizen science that empower individuals to engage in data collection, analysis, and social work initiatives (Sultan et al., 2022). In this paradigm, particularly when server-assigned task mode (Chen, 2019) is followed, designing an effective task assignment and retention strategy is essential to sustainably maintain the volunteer workforce. This can be achieved by valuing their contributions through reimbursement or recognition and by modeling their latent potential for future goals (Kaur et al., 2022). Moreover, in case of complex and skill-demanding heterogeneous tasks, assignment become even more challenging because in that case candidates need to be chosen as per task’s explicit competency and skill requirements (Samanta et al., 2021; Cheng et al., 2016; Song et al., 2020; Hettiachchi et al., 2022). Thus, during retention, a proper balance among newcomers and experienced ones needs to be evaluated without discouraging any dropouts. In general, with respect to skill-based VCS ventures (Samanta et al., 2022), volunteers assigned to complex, heterogeneous and skill-oriented CS tasks, primarily participate for moral values and gaining experience. Thus, task assignment and retention strategies should focus on providing a positive impact on the participants’ satisfaction levels (Hettiachchi et al., 2022). Beyond remuneration support, these strategies should be driven by the experience gained and the skills developed by the participants.
@@
Figure 1. Food Bank: Skill-based Volunteer Crowdsourcing Service (VCS) Requiring Workforce Composition Balance for Regular Operations and Surge Events (Holidays, Natural Calamities, etc.)
1.1. Motivating Example
@@
Consider the example of a Food Bank run by an NGO (see Figure 1), which serves as a standard model in the realm of skill-based VCS, relying on volunteers for both long-term and short-term initiatives (Kaur et al., 2022; Wang et al., 2019). Skill demand is a critical aspect in this context, as volunteers are tasked with various responsibilities such as sorting and packaging food, cooking, auditing, driving delivery vehicles, and managing logistics. Apart from addressing surge events like the distribution of food packages during specific high-demand periods (such as holidays, school breaks or natural calamities), the food bank requires a steady stream of volunteers to maintain long-term services. These services are supposed to operate on a regular schedule, such as opening for fixed hours on 2–3 designated days each week. While new volunteers are always welcomed, it is equally important to properly incentivize existing volunteers based on their latent potential and expected future contributions.
2. Limitations, Challenges and Our Research Questions
@@
    Figure 2. Proposed Worker Composition Balance (WCB) Framework: Integrating Task Assignment Mechanism with Volunteer Retention and Value Enhancement (VRAVE) Module
2.1. Contributions
In this paper, the aforementioned research questions are addressed by focusing on developing a workforce composition balance policy framework aimed at retaining potential volunteers for sustained participation. The policy is specifically designed to seamlessly integrate with the existing skill-oriented task assignment strategy. For our methodology, we chose the state-of-the-art SWill-TAC (Samanta et al., 2021) as the foundational skill-oriented task assignment method. SWill-TAC, a skill-oriented dynamic task allocation approach with a willingness factor for complex assignments, iteratively assigns workers (volunteers) to tasks based on the required skills and their skill sets. SWill-TAC also considers the workers’ willingness to perform tasks and monitors the budget constraints associated with the tasks. In addition to our proposed retention mechanism, we conducted experiments with three baseline retention mechanisms proposed by the authors in (Difallah et al., 2014). All the experiments are conducted on a real life dataset i.e UpWork (Upwork, 2021). The main contributions of this paper are:
@@ 3. Related Work
The authors in (Pilourdault et al., 2017) attempted to model worker motivation by examining its effect on different performance metrics in micro-tasks CS. They allowed workers to select tasks and proposed three mechanisms: task relevance-based, task diversity-based, and combined diversity-payment-based. However, their focus was primarily on motivating participants who are already working on a batch of micro-tasks but are reluctant to contribute their full potential.
Similarly, (Difallah et al., 2014) aimed to reduce overall batch execution latency by retaining participants working on a particular batch of HITs, with experiments conducted on Amazon MTurk. This study concentrated on maintaining worker participation for the duration of specific tasks to improve efficiency.
In (Kaur et al., 2022), the authors created an integer program to help non-profit organizations generate volunteer and employee task assignments for specific program planning horizons. This model takes into account volunteer retention needs, as well as organizational and task urgency needs when assigning tasks.
@@ 4. Preliminaries
The proposed framework operates under the assumption that a task’s completion depends on the coverage of skills required for that task. However, volunteers differ from regular paid crowd workers in that they receive a remuneration $RW_{v}$ specifically to cover their incurred expense $C_{v}$. Moreover, it is important to note that $RW_{v}$ is not always equal to $C_{v}$. We also assume that $C_{v}$ remains the same for any given task assigned to volunteer $v$. Similar to the approaches taken by authors in (Song et al., 2020; Samanta et al., 2021, 2022; Cheng et al., 2016), we consider that the skills required for a task or possessed by a volunteer are always part of a preset universal skill set $S$ with a constant size. Additionally, following the concept of one-to-many mapping as discussed in (Samanta et al., 2021), it is possible for a task to have multiple volunteers at a given time frame, but the reverse scenario is not true.
@@
(2) $\Pi_{init}(v,r)=\frac{1}{1+\exp^{(-\sigma_{(v,r)})}}$
@@
(4) $\text{And}\>\>\psi_{v}=\alpha^{1/l_{v}}$
To begin, $\sigma_{(v,r)}$ comprises three additive components. The first component is the ratio of the number of times a volunteer $v$ has been successfully selected ($Alloc^{succ}_{v}$) to the total number of allocation rounds $v$ has participated in so far ($Alloc^{par}_{v}$). This ratio helps gauge the historical performance of volunteers. The second component is the ratio of the skills possessed by $v$ to the total number of skills in the universal skill set $S$. This ratio reflects the current competencies of volunteers given the diverse task requirements in the pool. Many CS platforms, such as Upwork(Upwork, 2021), provide skill development resources through which crowd-workers can update and enhance their skill sets. Therefore, our framework assumes that over execution rounds, the skill set $S_{v}$ of volunteers will expand.
@@
The Aging Factor incorporates the aging constant $\alpha$, which ranges from $(0,1]$. The input $l$ represents the number of rounds that have passed since a volunteer $v$ was last successfully assigned to a task. As $v$ remains unallocated for more rounds, $\psi_{v}$ increases. For newcomers, the aging constant is set to $0.1$. The concepts of $\psi$ and $\Pi_{init}$ are inspired by operating system principles and related works (Silberschatz et al., 2006; Lindgren et al., 2004; Doria and Lindgren, 2015).
@@
(5) $D_{v}=\gamma\cdot\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)\cdot\left(\frac{U}{\sum_{v\in{V_{unassigned}}}\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)}\right)$
@@
(6) $\text{Sat}(v,r)=(1-\omega).\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)+\omega.\left(\frac{CI(v,r)}{C_{v}\times(r-1)}\right)$
@@
(7) $CI(v,r)=CI(v,r-1)+C_{v}+D_{v}$
5. Workforce Composition Balance Framework
In this section, we build upon the preliminary definitions presented earlier to delve into the workflow of our proposed Workforce Composition Balance (WCB) Framework. We initiate by outlining the Volunteer Retention and Value Enhancement (VRAVE) algorithm (refer to Algorithm 1), which addresses research question RQ2. Following that, we explore the seamless integration of VRAVE with a skill-oriented task assignment method tailored for managing complex heterogeneous tasks. Specifically, we highlight SWill-TAC (Samanta et al., 2021) as our chosen foundational skill-oriented task assignment method within this framework. Therefore, this integration feature of the WCB framework provides the answer to research question RQ1. It is worth noting that the workflow of our framework can also accommodate other skill-oriented task assignment method in place of SWill-TAC.
@@ 6. Evaluation
6.1. Datasets Description
We used the Upwork dataset mentioned in (Samanta et al., 2021, 2022), following a common practice adopted by authors in (Jiang et al., 2020; Fu et al., 2015; Yadav et al., 2021; Samanta et al., 2021, 2022). UpWork is a marketplace that leverages a trust-based network linking clients and freelancers. The website provides information about available crowd workers (in our case, volunteers) and jobs that have been posted. We used the tags associated with crowd workers’ profiles to determine their available skills. The tasks list the required skills, time, and budget. During data pre-processing, we eliminated any duplicate tasks or crowd workers to reduce redundancy. The attributes included in the task category are task name, ID, budget amount, expected completion time in hours, skill specifications, and arrival time. For the crowd worker category, the attributes are crowd worker name, ID, efficiency, skills, demanded payment (will be replicated as incurred expense), and arrival time. The bias factor and efficiency are calculated using a Uniform distribution, $U(0,1)$ (which would be used by SWill-TAC). Both the task and worker arrival sequences are assumed to follow a Poisson distribution. Since the number of workers is about 15 times larger than the number of tasks, the rate of arrival for workers is estimated at $75$ per unit of time, and the rate for tasks is $5$ per unit of time. The parameter settings used in our experiments are detailed in Table-1.
Table 1. Parameter settings for experiment
@@ 6.2. Baselines
One such state-of-the-art work is presented by Difallah et al. in (Difallah et al., 2014). In their study, they introduce pricing schemes aimed at enhancing the retention rate of workers engaged in long batches of similar tasks. They investigate how varying the monetary reward over time affects the willingness of workers to complete tasks within a batch. Their work compares new pricing schemes against traditional methods, such as offering a constant reward for all tasks in a batch, and empirically demonstrates how certain schemes effectively incentivize workers to continue working on a given batch of tasks for an extended period. Consequently, we developed three baseline models by integrating the assignment logic of SWill-TAC with the incentive policies from (Difallah et al., 2014).
@@ 6.3. Results and Discussion
Figure 3. Average satisfaction scores over six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
@@
Figure 4. Total completed tasks and retained volunteers per six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
@@
Figure 5. Average remuneration paid over six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
7. Conclusion and Future Work

### v2

@@ Abstract
 [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
Crowdsourcing (CS) has become integral to gathering and analyzing diverse data for various social applications like volunteering, delivery logistics, and participatory sensing [1]. These activities involve diverse crowd workers who contribute actively or passively, dealing with spatial or non-spatial data in various formats and qualities.
@@ 1 Introduction
Research shows that most of the crowd workers on online CS platforms are primarily short-term participants. This deficiency in sustained crowd engagement has led to an imbalance in workforce composition, hindering large market enterprises from achieving their full utility [2]. Crowd workers often disengage due to waning motivation or switch platforms for better perks and visibility, highlighting the need for strategies to select and retain competent workers [3]. While initial participation is vital in CS platforms, sustained crowd engagement is crucial for their long-term success and the continuous delivery of high-quality services for social good [4, 5].
@@
Volunteer Crowdsourcing Services (VCS), such as VolunteerMatch [6] and SkillImpact [7], are platforms within the growing community of citizen science that empower individuals to engage in data collection, analysis, and social work initiatives [8]. In this paradigm, particularly when server-assigned task mode [9] is followed, designing an effective task assignment and retention strategy is essential to sustainably maintain the volunteer workforce. This can be achieved by valuing their contributions through reimbursement or recognition and by modeling their latent potential for future goals [10]. Moreover, in case of complex and skill-demanding heterogeneous tasks, assignment become even more challenging because in that case candidates need to be chosen as per task’s explicit competency and skill requirements [11, 12, 13, 14]. Thus, during retention, a proper balance among newcomers and experienced ones needs to be evaluated without discouraging any dropouts. In general, with respect to skill-based VCS ventures [15], volunteers assigned to complex, heterogeneous and skill-oriented CS tasks, primarily participate for moral values and gaining experience. Thus, task assignment and retention strategies should focus on providing a positive impact on the participants’ satisfaction levels [14]. Beyond remuneration support, these strategies should be driven by the experience gained and the skills developed by the participants.
@@
Figure 1: Food Bank: Skill-based Volunteer Crowdsourcing Service (VCS) Requiring Workforce Composition Balance for Regular Operations and Surge Events (Holidays, Natural Calamities, etc.)
1.1 Motivating Example
@@
Consider the example of a Food Bank run by an NGO (see Figure 1), which serves as a standard model in the realm of skill-based VCS, relying on volunteers for both long-term and short-term initiatives [10, 16]. Skill demand is a critical aspect in this context, as volunteers are tasked with various responsibilities such as sorting and packaging food, cooking, auditing, driving delivery vehicles, and managing logistics. Apart from addressing surge events like the distribution of food packages during specific high-demand periods (such as holidays, school breaks or natural calamities), the food bank requires a steady stream of volunteers to maintain long-term services. These services are supposed to operate on a regular schedule, such as opening for fixed hours on 2–3 designated days each week. While new volunteers are always welcomed, it is equally important to properly incentivize existing volunteers based on their latent potential and expected future contributions.
2 Limitations, Challenges and Our Research Questions
@@
    Figure 2: Proposed Worker Composition Balance (WCB) Framework: Integrating Task Assignment Mechanism with Volunteer Retention and Value Enhancement (VRAVE) Module
2.1 Contributions
In this paper, the aforementioned research questions are addressed by focusing on developing a workforce composition balance policy framework aimed at retaining potential volunteers for sustained participation. The policy is specifically designed to seamlessly integrate with the existing skill-oriented task assignment strategy. For our methodology, we chose the state-of-the-art SWill-TAC [11] as the foundational skill-oriented task assignment method. SWill-TAC, a skill-oriented dynamic task allocation approach with a willingness factor for complex assignments, iteratively assigns workers (volunteers) to tasks based on the required skills and their skill sets. SWill-TAC also considers the workers’ willingness to perform tasks and monitors the budget constraints associated with the tasks. In addition to our proposed retention mechanism, we conducted experiments with three baseline retention mechanisms proposed by the authors in [17]. All the experiments are conducted on a real life dataset i.e UpWork [18]. The main contributions of this paper are:
@@ 3 Related Work
The authors in [19] attempted to model worker motivation by examining its effect on different performance metrics in micro-tasks CS. They allowed workers to select tasks and proposed three mechanisms: task relevance-based, task diversity-based, and combined diversity-payment-based. However, their focus was primarily on motivating participants who are already working on a batch of micro-tasks but are reluctant to contribute their full potential.
Similarly, [17] aimed to reduce overall batch execution latency by retaining participants working on a particular batch of HITs, with experiments conducted on Amazon MTurk. This study concentrated on maintaining worker participation for the duration of specific tasks to improve efficiency.
In [10], the authors created an integer program to help non-profit organizations generate volunteer and employee task assignments for specific program planning horizons. This model takes into account volunteer retention needs, as well as organizational and task urgency needs when assigning tasks.
@@ 4 Preliminaries
The proposed framework operates under the assumption that a task’s completion depends on the coverage of skills required for that task. However, volunteers differ from regular paid crowd workers in that they receive a remuneration $RW_{v}$ specifically to cover their incurred expense $C_{v}$. Moreover, it is important to note that $RW_{v}$ is not always equal to $C_{v}$. We also assume that $C_{v}$ remains the same for any given task assigned to volunteer $v$. Similar to the approaches taken by authors in [13, 11, 15, 12], we consider that the skills required for a task or possessed by a volunteer are always part of a preset universal skill set $S$ with a constant size. Additionally, following the concept of one-to-many mapping as discussed in [11], it is possible for a task to have multiple volunteers at a given time frame, but the reverse scenario is not true.
@@
$\Pi_{init}(v,r)=\frac{1}{1+\exp^{(-\sigma_{(v,r)})}}$ (2)
@@
$\text{And}\>\>\psi_{v}=\alpha^{1/l_{v}}$ (4)
To begin, $\sigma_{(v,r)}$ comprises three additive components. The first component is the ratio of the number of times a volunteer $v$ has been successfully selected ($Alloc^{succ}_{v}$) to the total number of allocation rounds $v$ has participated in so far ($Alloc^{par}_{v}$). This ratio helps gauge the historical performance of volunteers. The second component is the ratio of the skills possessed by $v$ to the total number of skills in the universal skill set $S$. This ratio reflects the current competencies of volunteers given the diverse task requirements in the pool. Many CS platforms, such as Upwork[18], provide skill development resources through which crowd-workers can update and enhance their skill sets. Therefore, our framework assumes that over execution rounds, the skill set $S_{v}$ of volunteers will expand.
@@
The Aging Factor incorporates the aging constant $\alpha$, which ranges from $(0,1]$. The input $l$ represents the number of rounds that have passed since a volunteer $v$ was last successfully assigned to a task. As $v$ remains unallocated for more rounds, $\psi_{v}$ increases. For newcomers, the aging constant is set to $0.1$. The concepts of $\psi$ and $\Pi_{init}$ are inspired by operating system principles and related works [20, 21, 22].
@@
$\begin{split}D_{v}=\gamma\cdot\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)&\cdot\\ \left(\frac{U}{\sum_{v\in{V_{unassigned}}}\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)}\right)\end{split}$ (5)
@@
$\text{Sat}(v,r)=(1-\omega).\left(\frac{\Pi(v,r-1)}{\max_{\forall v^{\prime}\in V\text{ at }r}\Pi(v^{\prime},r-1)}\right)+\omega.\left(\frac{CI(v,r)}{C_{v}\times(r-1)}\right)$ (6)
@@
$CI(v,r)=CI(v,r-1)+C_{v}+D_{v}$ (7)
5 Workforce Composition Balance Framework
In this section, we build upon the preliminary definitions presented earlier to delve into the workflow of our proposed Workforce Composition Balance (WCB) Framework. We initiate by outlining the Volunteer Retention and Value Enhancement (VRAVE) algorithm (refer to Algorithm 1), which addresses research question RQ2. Following that, we explore the seamless integration of VRAVE with a skill-oriented task assignment method tailored for managing complex heterogeneous tasks. Specifically, we highlight SWill-TAC [11] as our chosen foundational skill-oriented task assignment method within this framework. Therefore, this integration feature of the WCB framework provides the answer to research question RQ1. It is worth noting that the workflow of our framework can also accommodate other skill-oriented task assignment method in place of SWill-TAC.
@@ 6 Evaluation
6.1 Datasets Description
We used the Upwork dataset mentioned in [11, 15], following a common practice adopted by authors in [23, 24, 25, 11, 15]. UpWork is a marketplace that leverages a trust-based network linking clients and freelancers. The website provides information about available crowd workers (in our case, volunteers) and jobs that have been posted. We used the tags associated with crowd workers’ profiles to determine their available skills. The tasks list the required skills, time, and budget. During data pre-processing, we eliminated any duplicate tasks or crowd workers to reduce redundancy. The attributes included in the task category are task name, ID, budget amount, expected completion time in hours, skill specifications, and arrival time. For the crowd worker category, the attributes are crowd worker name, ID, efficiency, skills, demanded payment (will be replicated as incurred expense), and arrival time. The bias factor and efficiency are calculated using a Uniform distribution, $U(0,1)$ (which would be used by SWill-TAC). Both the task and worker arrival sequences are assumed to follow a Poisson distribution. Since the number of workers is about 15 times larger than the number of tasks, the rate of arrival for workers is estimated at $75$ per unit of time, and the rate for tasks is $5$ per unit of time. The parameter settings used in our experiments are detailed in Table-I.
TABLE I: Parameter settings for experiment
@@ 6.2 Baselines
One such state-of-the-art work is presented by Difallah et al. in [17]. In their study, they introduce pricing schemes aimed at enhancing the retention rate of workers engaged in long batches of similar tasks. They investigate how varying the monetary reward over time affects the willingness of workers to complete tasks within a batch. Their work compares new pricing schemes against traditional methods, such as offering a constant reward for all tasks in a batch, and empirically demonstrates how certain schemes effectively incentivize workers to continue working on a given batch of tasks for an extended period. Consequently, we developed three baseline models by integrating the assignment logic of SWill-TAC with the incentive policies from [17].
@@ 6.3 Results and Discussion
Figure 3: Average satisfaction scores over six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
@@
Figure 4: Total completed tasks and retained volunteers per six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
@@
Figure 5: Average remuneration paid over six execution rounds. The comparison is between the proposed VRAVE and three baselines, with SWill-TAC used as the task assignment mechanism.
7 Conclusion and Future Work